OpenStudy (anonymous):

If y varies directly as x and y is 180 when x is n and y is n when x is 5 what is the value of n

OpenStudy (dan815):

y=k*x 180 = k*n n=k*5 solve for n

OpenStudy (anonymous):

If f(×)=×-3÷× and g (×)=5×-4, what is the domain

OpenStudy (dan815):

do the first question

OpenStudy (anonymous):

N=5k

OpenStudy (anonymous):

Ok I got 180=5ksquared

OpenStudy (anonymous):

180=25k n=36

OpenStudy (anonymous):

Is n=36?